For a certain first order reaction, half life is 100s, how long will it take to be 75% completed

1
Expert's answer
2022-03-24T04:34:02-0400

t1/2 = 100 s;

t1/2 = ln(2)/k;

k = ln(2)/t1/2 = 0.693/100 = 0.00693 s-1;

[Ct]/[C0] = e-k*t;

ln([Ct]/[C0]) = -k*t;

k = -(ln([Ct]/[C0]))/t;

t = -(ln([Ct]/[C0]))/k = ln([C0]/[Ct])/k = ln(x/(x-0.75x))/k = ln(x/0.25x)/k = ln(4)/0.00693 = 200 s.

Answer: 200 s.